OpenStudy (anonymous):

An amount of $41,000 is borrowed for 14 years at 4.5% interest, compounded annually. If the loan is paid in full at the end of that period, how much must be paid back? Round your answer to the nearest dollar.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

using the general formula for compounded interest: A=P(1+r/n)^(nt) A=41000(1+0.045/1)^(1*14) A= $75929.74